There are 8 action, 3 comedy, and 5 children's DVDs on a shelf. Suppose two DVDs are selected at random from the shelf. Find each probability
P(2 action DVD) if no replacement occurs
P(a comedy DVD, then children's DVD) if no replacement occurs

Answer by edjones(8007)

There is a total of 16 DVDs.
P(2 action DVD) if no replacement occurs
8/16 * 7/15 = 7/30
.
P(a comedy DVD, then children's DVD) if no replacement occurs
3/16 * 5/15 = 1/16
.
Ed